OpenStudy (anonymous):

Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid $135. This year, he paid $162 . What was the percent increase in the cost of his plane ticket? Show your work.

OpenStudy (anonymous):

The increase is taken as current rate-last year's rate=162-135=27$ To take the percentage you take= (increase/past year's rate)*100 = 27/135 *100 = 20%